x-9y=46 7y-10x=-45 I"m substituting

Answer by dfrazzetto(283)

x-9y=46 and 7y-10x=-45, using substitution:
x=9y+46, substitute into second equation:
7y - 10(9y+46) = -45
7y -90y -460 = -45
-83y - 460 = -45
-83y = 415
y = -5
put y=5 back into one of the original equations:
x=9y + 46
x=9(-5) +46
x = -45 + 46
x=1
CHECK:
x=1, y=-5
7y-10x = -45
7(-5) - 10(1) = -45
-35 - 10 = -45
-45=-45 √